OpenStudy (anonymous):

Solve the equation 70 = –7(–2 – 2z) A)z = 4 B)z = –28 C)z = –112 D)z = 784

OpenStudy (anonymous):

Hello Breannya. 70 = -7(-2-2z) -10 = -2 - 2z -8 = -2z z = 4 So it is A

OpenStudy (saifoo.khan):

Rightt.